OpenStudy (dls):

If the mass of earth is assumed to be m,the radius to be r,the angular velocity be w(wrt centre) and w' wrt sun ,the distance between the centre of earth and sun be l,what will be the earth's kinetic energy?

OpenStudy (dls):

@Mashy @RnR rotational dynamics :D

OpenStudy (anonymous):

shimple :D

OpenStudy (anonymous):

add the two kinetic energies.. K.E(rotation) + K.E(revolution)

OpenStudy (dls):

:O

OpenStudy (anonymous):

energies are scalar.. they can be directly added !

OpenStudy (dls):

how do i calculate them? :(

OpenStudy (anonymous):

what will be K.E(rotation) ?

OpenStudy (dls):

\[\LARGE \frac{1}{2}I \omega ^2\]

OpenStudy (dls):

^___________________________________^

OpenStudy (anonymous):

what is I when you consider earths rotation!?

OpenStudy (dls):

Mr^2?

OpenStudy (anonymous):

DON"T YOU KNOW THE MOMENT OF INERTIA OF A SPHERE ABOUT ITS CENTRE?!??!?!?!!?!?

OpenStudy (dls):

its a hollow sphere! so 2/4MR^2

OpenStudy (anonymous):

:O.. its a solid sphere .. !!!

OpenStudy (dls):

oh wth earth is nothollow lol 2/5mr^2 :P

OpenStudy (anonymous):

LOL

OpenStudy (dls):

lol nevermind

OpenStudy (anonymous):

wow come to think of it.. totally sidelining i had thought moment of inertia of a hollow sphere of mass M and radius R would be MR^2 itself.. cause every particle is at the same distance R from the centre.. why is it 2/3 (MR^2) :O

OpenStudy (anonymous):

oh no never mind :D i got it!!

OpenStudy (dls):

:/ great

OpenStudy (anonymous):

so anwasy .. now come back so what is the K.E of rotation!?

OpenStudy (dls):

\[\LARGE \frac{1}{2} \frac{2}{3}M R^2 \omega\]

OpenStudy (anonymous):

correct now when you calculate the K.E of revolution.. what is the moment of inerita?

OpenStudy (anonymous):

2/3 or 2/5? :O

OpenStudy (anonymous):

and wait fool.. its not a hollow sphere.. its a SOLID sphere.. sorry.. i dunno M.I values..

OpenStudy (dls):

we will calculate like this?  ML^2 ? wahi to?:/ 2/5 hoga

OpenStudy (anonymous):

NO.. this time is a point mass (earth).. spinning about sun.. so it ll be just ML^2